# Benefits of Early Stopping in Gradient Descent for Overparameterized Logistic Regression

### v1

@@ 1 Introduction
The implicit regularization of GD is relatively well understood in regression settings. Using overparameterized linear regression as an example, amongst all interpolators, GD asymptotically converges to the minimum $\ell_{2}$-norm interpolator (Zhang et al., 2021). Moreover, when the data covariance satisfies certain conditions, the minimum $\ell_{2}$-norm interpolator achieves vanishing excess risk while fitting training data with constant amount of noise, a phenomenon known as benign overfitting (see Bartlett et al., 2020; Tsigler and Bartlett, 2023, and references therein). When the data covariance is general, although benign overfitting may not occur, early-stopped GD (and one-pass stochastic GD) can still achieve vanishing excess risk (Bühlmann and Yu, 2003; Yao et al., 2007; Lin and Rosasco, 2017; Dieuleveut and Bach, 2016; Zou et al., 2023, 2022; Wu et al., 2022a). This suggests early stopping provides an additional regularization effect for GD in linear regression. Moreover, the statistical effects of early stopping are known to be comparable to that of $\ell_{2}$-regularization in linear regression (Suggala et al., 2018; Ali et al., 2019; Zou et al., 2021).
However, the picture is less complete for classification, where the risk is measured by the logistic loss and the zero-one loss instead of the squared loss. In overparameterized logistic regression, GD diverges in norm while converging in direction to the maximum $\ell_{2}$-margin solution (see Soudry et al., 2018; Ji and Telgarsky, 2018, and Proposition 2.2 in Section 2), which is in contrast with GD’s convergence to the (bounded!) minimum $\ell_{2}$-norm solution in the linear regression setting. In standard (finite-dimensional, low-noise, large margin) classification settings, the asymptotic implicit bias of GD implies generalization via classical margin theory (Bartlett and Shawe-Taylor, 1999). More recently, certain high-dimensional settings exhibit well-behaved maximum margin solutions and benign overfitting (see Montanari et al., 2019, for example), but it is unclear if these results apply more broadly or represent special cases. Moreover, if the maximum $\ell_{2}$-margin solution generalizes poorly, new techniques are required, as the aforementioned least squares techniques cannot be easily adapted owing to their heavy dependence upon the explicit linear algebraic form of GD’s path specific to least squares.
@@ 2 Preliminaries
We say an estimator $\hat{\mathbf{w}}$ is consistent (for classification) if it attains the Bayes zero-one error asymptotically, that is, $\mathcal{E}(\hat{\mathbf{w}})-\min\mathcal{E}\to 0$. We say an estimator $\hat{\mathbf{w}}$ is calibrated if its induced conditional probability predicts the true one asymptotically (Foster and Vohra, 1998), that is, $\mathcal{C}(\hat{\mathbf{w}})\to 0$.
@@
The first condition ensures the population distribution carries a constant amount of noise, as the Bayes logistic risk and Bayes zero-one error are lower bounded by a constant according to Proposition 2.1. In other words, the population distribution is strictly not linearly separable. Despite so, the second condition ensures the linear separability of the training data almost surely, as the number of effective parameters exceeds the number of training data. In this regime, estimators can interpolate the training data, yet this interpolation inherently carries the risk of overfitting and poor calibration. Our setting aligns well with the prior setting for studying benign overfitting in linear regression (Bartlett et al., 2020; Tsigler and Bartlett, 2023).
@@
When the training data is linearly separable (implied by overparameterization), prior works show that GD diverges to infinite in norm while converging in direction to the maximum $\ell_{2}$-margin direction (Soudry et al., 2018; Ji and Telgarsky, 2018). This characterizes the asymptotic implicit bias of GD. See the following proposition for a precise statement.
@@
For logistic regression with linearly separable data, the implicit bias of GD is established by Soudry et al. (2018); Ji and Telgarsky (2018) when the stepsizes are small such that ^$\widehat{\mathcal{L}}(\mathbf{w}_{t})$ decreases monotonically. The same results can be extended to GD with any fixed stepsize using techniques from (Wu et al., 2023, 2024). ∎
@@ 3.1 A Bias-Dominating Bound
where $C_{0}>1$ is a universal constant. Then with probability at least $1-\delta$, the empirical risk ^$\widehat{\mathcal{L}}(\mathbf{w}_{t})$ decreases monotonically for $t\geq 0$. Thus there exists a stopping time $t$ such that
@@
${\mathcal{L}(\mathbf{w}_{t})-\min\mathcal{L}\lesssim}\sqrt{\frac{\max\big{\{}1,\,\operatorname{tr}(\bm{\Sigma})\|\mathbf{w}^{*}_{0:k}\|^{2}\big{\}}\ln^{2}(n/\delta)}{n}}+\|\mathbf{w}^{*}_{k:\infty}\|^{2}_{\bm{\Sigma}}.$
A few remarks on Theorem 3.1 are in order.
@@
As a concrete example, let us consider the following source and capacity conditions (Caponnetto and De Vito, 2007),
@@ 3.2 A Variance-Dominating Bound
$\mathcal{L}(\mathbf{w}_{t})-\min\mathcal{L}\lesssim\|\mathbf{w}^{*}\|\bigg{(}\frac{k}{n}+\sqrt{\frac{\sum_{i>k}\lambda_{i}}{n}}+\frac{\operatorname{tr}(\bm{\Sigma})^{1/2}\ln\big{(}n\|\mathbf{w}^{*}\|\operatorname{tr}(\bm{\Sigma})/\delta\big{)}}{n}\bigg{)}.$
@@
Theorems 3.1 and 3.2 are sufficiently powerful for our purpose of demonstrating the benefits of early stopping. However, we point out that neither Theorems 3.1 nor 3.2 reveal the true trade-off between the bias and variance errors induced by early stopping. This is unsatisfactory given that in linear regression, the exact trade-off between bias and variance errors has been settled for one-pass SGD (Zou et al., 2023; Wu et al., 2022a, b) and $\ell_{2}$-regularization (Tsigler and Bartlett, 2023), and has been partially settled for early-stopped GD (Zou et al., 2022, assuming a Gaussian prior). We leave the improvement of these bounds for future work.
@@ 3.3 Implicit Bias of Early Stopping
In this part, we briefly discuss the proof ideas by introducing the following key lemma in our analysis. Variants of this lemma have appeared in (Ji and Telgarsky, 2018, 2019; Shamir, 2021; Telgarsky, 2022; Wu et al., 2024) for analyzing different aspects of GD. For completeness, we include a proof of it in Section A.2.
@@
Besides Lemma 3.3, the remaining efforts for proving Theorems 3.1 and 3.2 are using classical tools of Rademacher complexity (Bartlett and Mendelson, 2002; Kakade et al., 2008) and local Rademacher complexity (Bartlett et al., 2005), respectively. More details can be found in Sections B.1 and B.2.
@@ 4.2 Zero-One Error
Theorem 4.2 characterizes a class of overparameterized logistic regression problems where every interpolating estimator needs at least an exponential number of training data to achieve a small excess zero-one error. This applies to asymptotic GD as it converges to the maximum $\ell_{2}$-margin solution by Proposition 2.2. In contrast, Theorems 3.1 and 3.2 suggests that early-stopped GD can achieve a small excess zero-one error using at most a polynomial number of training data under weak conditions. These weak conditions can be, for example, $\|\mathbf{w}^{*}\|<\infty$ or the sparsity parameter $k$ does not grow with $n$. This separation underscores the benefits of early stopping for reducing sample complexity.
@@ 5 Early Stopping and $\ell_{2}$-Regularization
In this section, we establish connections between early stopping and $\ell_{2}$-regularization. An $\ell_{2}$-regularized empirical risk minimizer (ERM) is defined as
@@ 5.1 A Global Connection
$\cos(\mathbf{w}_{t},\mathbf{u}_{\lambda})\geq\frac{1}{\sqrt{2}},\quad\frac{\sqrt{2}}{1+\sqrt{2}}\|\mathbf{u}_{\lambda}\|\leq\|\mathbf{w}_{t}\|\leq\frac{\sqrt{2}}{\sqrt{2}-1}\|\mathbf{u}_{\lambda}\|.$
@@ 5.2 An Asymptotic Comparison
In logistic regression with linearly separable data, both GD and $\ell_{2}$-regularization paths diverge to infinite in norm (as $t\to\infty$ and $\lambda\to 0$) while converging in direction to the maximum $\ell_{2}$-margin solution (Rosset et al., 2004; Soudry et al., 2018; Ji and Telgarsky, 2018; Ji et al., 2020). Therefore their angle tends to zero asymptotically (Suggala et al., 2018; Ji et al., 2020). This characterization is more precise than the $\pi/4$ global angle bound from Theorem 5.1.
However, it remains unclear how the $\ell_{2}$-distance between the two paths evolves in logistic regression with linearly separable data. Quite surprisingly, we will show that their $\ell_{2}$-distance tends to zero under a widely used condition (Soudry et al., 2018; Ji and Telgarsky, 2021; Wu et al., 2023), but could diverge to infinity in the worst case.
@@
Here, $\bm{\beta}$ are the dual variables multiplied by $\mathbf{y}$ entry-wise. Let $\hat{\bm{\beta}}$ be the solution to the above problem. Let $\mathcal{S}:=\{i\in[n]:y_{i}\hat{\bm{\beta}}_{i}>0\}$ be the set of support vectors (with strictly positive dual variables). The following condition assumes the coverage of the support vectors.
@@
Assume that $\operatorname{rank}\{\mathbf{x}_{i}:i\in\mathcal{S}\}=\operatorname{rank}\{\mathbf{x}_{i}:i\in[n]\}$.
Assumption 2 has been widely used in the analysis of the implicit bias (Soudry et al., 2018; Ji and Telgarsky, 2021; Wu et al., 2023). In particular, Assumption 2 holds if every data is a support vector, which is common in high-dimensional situations (Hsu et al., 2021; Wang and Thrampoulidis, 2022; Cao et al., 2021).
@@
We conjecture that for logistic regression with linearly separable data, the limit of the $\ell_{2}$-distance between the GD and $\ell_{2}$-regularized paths is either zero or infinity, and the phase transition is determined by a certain geometric property of the dataset (for example, Assumption 2). This is left for future investigation.
@@ 6 Related Works
A line of work shows the benign overfitting of the asymptotic GD (or the maximum $\ell_{2}$-margin estimator) in overparameterized logistic regression under a variety of assumptions (Montanari et al., 2019; Chatterji and Long, 2021; Cao et al., 2021; Wang and Thrampoulidis, 2022; Muthukumar et al., 2021; Shamir, 2023). Our results are not a violation of theirs, instead, we show an additional regularization of early-stopping, which brings statistical advantages of early-stopped GD over asymptotic GD such as calibration and a smaller sample complexity.
@@
In the classical finite $d$-dimensional setting, the sample complexity of the empirical risk minimizer (ERM) for logistic regression is well-studied (Ostrovskii and Bach, 2021; Kuchelmeister and van de Geer, 2024; Hsu and Mazumdar, 2024; Chardon et al., 2024), where the minimax rate is known to be $\mathcal{O}(d/n)$. Different from theirs, we focus on an overparameterized regime, where the ERM of logistic regression does not even exist. When specialized to their setting, our Theorem 3.2 recovers the comparable $\tilde{\mathcal{O}}(d/n)$ rate.
@@
Finally, Bach (2014) considered one-pass SGD for logistic regression assuming strong convexity around the true model parameter. This strong convexity assumption, however, is prohibitive in our high-dimensional settings.
@@
There are logistic risk bounds of early-stopped GD (and one-pass SGD) developed in the noiseless cases, assuming a separable population distribution (Ji and Telgarsky, 2018; Shamir, 2021; Telgarsky, 2022; Schliserman and Koren, 2024). These results do not imply any benefits of early stopping, as their setting is noiseless. In comparison, we consider overparameterized logistic regression with a strictly non-separable population distribution, where the risk of overfitting is prominent. In this case, our results suggest that early stopping plays a significant role in preventing overfitting.
@@
In the boosting literature, an early work by Zhang and Yu (2005) showed that boosting methods (that can be interpreted as coordinate descent) with early stopping are consistent in the classification sense; related refined studies for boosting with the squared loss with early stopping were also provided by Bühlmann and Yu (2003). The paper is also notable for giving the first proof of boosting methods converging to the maximum margin solution (Zhang and Yu, 2005, Appendix D), which was later refined with rates by (Telgarsky, 2013). Their results can be converted to GD. In particular, related concepts were used to prove consistency of early-stopped GD for shallow networks in the lazy regime (Ji et al., 2021). In contrast with the present work that focuses on high-dimensional cases, the preceding works only deal with finite-dimensional settings. Moreover, none of those works provide lower bounds for interpolating estimators and tight links to the regularization path which are provided in the present work.

### v2

@@ Abstract
Implicit Bias, Gradient Descent, Early Stopping, Logistic Regression, Overparameterization
@@ 1 Introduction
The implicit regularization of GD is relatively well understood in regression settings. Using overparameterized linear regression as an example, amongst all interpolators, GD asymptotically converges to the minimum $\ell_{2}$-norm interpolator (Zhang et al., 2021). Moreover, when the data covariance satisfies certain conditions, the minimum $\ell_{2}$-norm interpolator achieves vanishing excess risk while fitting training data with constant amount of noise, a phenomenon known as benign overfitting (see Bartlett et al., 2020; Tsigler & Bartlett, 2023, and references therein). When the data covariance is general, although benign overfitting may not occur, early-stopped GD (and one-pass stochastic GD) can still achieve vanishing excess risk (Bühlmann & Yu, 2003; Yao et al., 2007; Lin & Rosasco, 2017; Dieuleveut & Bach, 2016; Zou et al., 2023, 2022; Wu et al., 2022a). This suggests early stopping provides an additional regularization effect for GD in linear regression. Moreover, the statistical effects of early stopping are known to be comparable to that of $\ell_{2}$-regularization in linear regression (Suggala et al., 2018; Ali et al., 2019; Zou et al., 2021; Sonthalia et al., 2024).
However, the picture is less complete for classification, where the risk is measured by the logistic loss and the zero-one loss instead of the squared loss. In overparameterized logistic regression, GD diverges in norm while converging in direction to the maximum $\ell_{2}$-margin solution (see Soudry et al., 2018; Ji & Telgarsky, 2018, and Proposition 2.2 in Section 2), which is in contrast with GD’s convergence to the (bounded!) minimum $\ell_{2}$-norm solution in the linear regression setting. In standard (finite-dimensional, low-noise, large margin) classification settings, the asymptotic implicit bias of GD implies generalization via classical margin theory (Bartlett & Shawe-Taylor, 1999). More recently, certain high-dimensional settings exhibit well-behaved maximum margin solutions and benign overfitting (see Montanari et al., 2019, for example), but it is unclear if these results apply more broadly or represent special cases. Moreover, if the maximum $\ell_{2}$-margin solution generalizes poorly, new techniques are required, as the aforementioned least squares techniques cannot be easily adapted owing to their heavy dependence upon the explicit linear algebraic form of GD’s path specific to least squares.
@@ 2 Preliminaries
We say an estimator $\hat{\mathbf{w}}$ is consistent (for classification) if it attains the Bayes zero-one error asymptotically, that is, $\mathcal{E}(\hat{\mathbf{w}})-\min\mathcal{E}\to 0$. We say an estimator $\hat{\mathbf{w}}$ is calibrated if its induced conditional probability predicts the true one asymptotically (Foster & Vohra, 1998), that is, $\mathcal{C}(\hat{\mathbf{w}})\to 0$.
@@
The first condition ensures the population distribution carries a constant amount of noise, as the Bayes logistic risk and Bayes zero-one error are lower bounded by a constant according to Proposition 2.1. In other words, the population distribution is strictly not linearly separable. Despite so, the second condition ensures the linear separability of the training data almost surely, as the number of effective parameters exceeds the number of training data. In this regime, estimators can interpolate the training data, yet this interpolation inherently carries the risk of overfitting and poor calibration. Our setting aligns well with the prior setting for studying benign overfitting in linear regression (Bartlett et al., 2020; Tsigler & Bartlett, 2023).
@@
When the training data is linearly separable (implied by overparameterization), prior works show that GD diverges to infinite in norm while converging in direction to the maximum $\ell_{2}$-margin direction (Soudry et al., 2018; Ji & Telgarsky, 2018). This characterizes the asymptotic implicit bias of GD. See the following proposition for a precise statement.
@@
For logistic regression with linearly separable data, the implicit bias of GD is established by Soudry et al. (2018); Ji & Telgarsky (2018) when the stepsizes are small such that ^$\widehat{\mathcal{L}}(\mathbf{w}_{t})$ decreases monotonically. The same results can be extended to GD with any fixed stepsize using techniques from (Wu et al., 2023, 2024). ∎
@@ 3.1 A Bias-Dominating Bound
where $C_{0}>1$ is a universal constant. Then with probability at least $1-\delta$, there exists a stopping time $t$ such that
@@
$\mathcal{L}(\mathbf{w}_{t})-\min\mathcal{L}\lesssim$
$\sqrt{\frac{\max\big{\{}1,\,\operatorname{tr}(\bm{\Sigma})\|\mathbf{w}^{*}_{0:k}\|^{2}\big{\}}\ln^{2}(n/\delta)}{n}}+\|\mathbf{w}^{*}_{k:\infty}\|^{2}_{\bm{\Sigma}}.$
The existence of the desired stopping time is because GD minimizes the empirical risk monotonically (Ji & Telgarsky, 2018). In Theorem 3.1, we choose $k$ to minimize the upper bounds. Intuitively, $k$ determines the number of dimensions in which early-stopped GD is able to learn the true parameter. Moreover, early-stopped GD ignores the remaining dimensions and pays an “approximation” error. A few more remarks on Theorem 3.1 are in order.
@@
As a concrete example, let us consider the following source and capacity conditions (Caponnetto & De Vito, 2007),
@@
Although the stopping time in Theorem 3.1 is implicit, one can compute an upper bound on it using standard optimization and concentration tools. Specifically, GD converges in $\mathcal{O}(1/t)$ rate as the empirical risk is convex and smooth. Moreover, we can compute ^$\widehat{\mathcal{L}}(\mathbf{w}_{0:k})$ using concentration bounds. These lead to an upper bound on the stopping time.
@@
We also note that the proof of Theorem 3.1 can be adapted to other loss functions that are convex, smooth, and Lipschitz.
@@ 3.2 A Variance-Dominating Bound
$\mathcal{L}(\mathbf{w}_{t})-\min\mathcal{L}\lesssim$ $\|\mathbf{w}^{*}\|\bigg{(}\frac{k}{n}+\sqrt{\frac{\sum_{i>k}\lambda_{i}}{n}}+$
$\qquad\frac{\operatorname{tr}(\bm{\Sigma})^{1/2}\ln\big{(}n\|\mathbf{w}^{*}\|\operatorname{tr}(\bm{\Sigma})/\delta\big{)}}{n}\bigg{)}.$
@@
Theorems 3.1 and 3.2 are sufficiently powerful for our purpose of demonstrating the benefits of early stopping. However, we point out that neither Theorems 3.1 nor 3.2 reveal the true trade-off between the bias and variance errors induced by early stopping. This is unsatisfactory given that in linear regression, the exact trade-off between bias and variance errors has been settled for one-pass SGD (Zou et al., 2023; Wu et al., 2022a, b) and $\ell_{2}$-regularization (Tsigler & Bartlett, 2023), and has been partially settled for early-stopped GD (Zou et al., 2022, assuming a Gaussian prior). We leave the improvement of these bounds for future work.
@@ 3.3 Implicit Bias of Early Stopping
In this part, we briefly discuss the proof ideas by introducing the following key lemma in our analysis. Variants of this lemma have appeared in (Ji & Telgarsky, 2018, 2019; Shamir, 2021; Telgarsky, 2022; Wu et al., 2024) for analyzing different aspects of GD. For completeness, we include a proof of it in Section A.2.
@@
Besides Lemma 3.3, the remaining efforts for proving Theorems 3.1 and 3.2 are using classical tools of Rademacher complexity (Bartlett & Mendelson, 2002; Kakade et al., 2008) and local Rademacher complexity (Bartlett et al., 2005), respectively. More details can be found in Sections B.1 and B.2.
@@ 4.2 Zero-One Error
Theorem 4.2 characterizes a class of overparameterized logistic regression problems where every interpolating estimator needs at least an exponential number of training data to achieve a small excess zero-one error. This applies to asymptotic GD as it converges to the maximum $\ell_{2}$-margin solution by Proposition 2.2. In contrast, Theorems 3.1 and 3.2 suggests that early-stopped GD can achieve a small excess zero-one error using at most a polynomial number of training data under weak conditions. These weak conditions can be, for example, $\|\mathbf{w}^{*}\|<\infty$ or the sparsity parameter $k$ does not grow with $n$ (see also the examples given by 1). This separation underscores the benefits of early stopping for reducing sample complexity.
The intuition behind Theorem 4.2 is that there are $k$ informative dimensions and a lot more uninformative dimensions. Since $n\gg k$, the training set cannot be separated purely using the $k$ informative dimensions. Thus, interpolators must use the uninformative dimensions to separate the data, leading to the risk lower bound.
@@ 5 Early Stopping and $\ell_{2}$-Regularization
Sections 3 and 4 demonstrate that early stopping carries a certain regularization effect that benefits its statistical performance. This regularization is, however, implicit. In this section, we attempt to provide some intuitions of the implicit regularization of early stopping by establishing its connections to an explicit, $\ell_{2}$-regularization. An $\ell_{2}$-regularized empirical risk minimizer (ERM) is defined as
@@ 5.1 A Global Connection
$\cos(\mathbf{w}_{t},\mathbf{u}_{\lambda})\geq\frac{1}{\sqrt{2}},$
$\frac{\sqrt{2}}{1+\sqrt{2}}\|\mathbf{u}_{\lambda}\|\leq\|\mathbf{w}_{t}\|\leq\frac{\sqrt{2}}{\sqrt{2}-1}\|\mathbf{u}_{\lambda}\|.$
@@ 5.2 An Asymptotic Comparison
In logistic regression with linearly separable data, both GD and $\ell_{2}$-regularization paths diverge to infinite in norm (as $t\to\infty$ and $\lambda\to 0$) while converging in direction to the maximum $\ell_{2}$-margin solution (Rosset et al., 2004; Soudry et al., 2018; Ji & Telgarsky, 2018; Ji et al., 2020). Therefore their angle tends to zero asymptotically (Suggala et al., 2018; Ji et al., 2020). This characterization is more precise than the $\pi/4$ global angle bound from Theorem 5.1.
However, it remains unclear how the $\ell_{2}$-distance between the two paths evolves in logistic regression with linearly separable data. Quite surprisingly, we will show that their $\ell_{2}$-distance tends to zero under a widely used condition (Soudry et al., 2018; Ji & Telgarsky, 2021; Wu et al., 2023), but could diverge to infinity in the worst case.
@@
Here, $\bm{\beta}$ are the dual variables multiplied by $\mathbf{y}$ entry-wise. Let $\hat{\bm{\beta}}$ be the solution to the above problem. Let $\mathcal{S}_{+}:=\{i\in[n]:y_{i}\hat{\bm{\beta}}_{i}>0\}$ be the set of support vectors (with strictly positive dual variables). The following condition assumes the coverage of the support vectors.
@@
Assume that $\operatorname{rank}\{\mathbf{x}_{i}:i\in\mathcal{S}_{+}\}=\operatorname{rank}\{\mathbf{x}_{i}:i\in[n]\}$.
Assumption 2 has been widely used in the analysis of the implicit bias (Soudry et al., 2018; Ji & Telgarsky, 2021; Wu et al., 2023). In particular, Assumption 2 holds if every data is a support vector, which is common in high-dimensional situations (Hsu et al., 2021; Wang & Thrampoulidis, 2022; Cao et al., 2021).
@@
We conjecture that for logistic regression with linearly separable data, the limit of the $\ell_{2}$-distance between the GD and $\ell_{2}$-regularized paths is either zero or infinity, and the phase transition is determined by a certain geometric property of the dataset (for example, Assumption 2). The reasoning behind this conjecture is as follows. Note that Assumption 2 implies that the dataset projected perpendicular to the max-margin direction (called “projected dataset”) is strictly nonseparable (Wu et al., 2023, Lemma 3.1). This is the main property used in Theorem 5.2. Moreover, in Theorem 5.3, the “projected dataset” is nonseparable but with margin zero—we conjecture this property is sufficient for Theorem 5.3 to hold. Now for a generic separable dataset, we check the “projected dataset”: if it is strictly nonseparable, Theorem 5.2 holds; if it is nonseparable but with margin zero, we conjecture Theorem 5.3 holds; otherwise, it is separable (with positive margin), we decompose the dataset recursively. This is the reasoning behind our conjecture.
It also remains unclear to what extent early stopping replicates the effects of explicit regularization for logistic regression. Specifically, is there a logistic regression example such that early-stopped GD has a better calibration/logistic risk rate than $\ell_{2}$-regularization or vice-versa? This is left for future investigation, as our current bounds are not sharp enough to yield a concrete answer.
@@ 6 Related Works
A line of work shows the benign overfitting of the asymptotic GD (or the maximum $\ell_{2}$-margin estimator) in overparameterized logistic regression under a variety of assumptions (Montanari et al., 2019; Chatterji & Long, 2021; Cao et al., 2021; Wang & Thrampoulidis, 2022; Muthukumar et al., 2021; Shamir, 2023). Our results are not a violation of theirs, instead, we show an additional regularization of early-stopping, which brings statistical advantages of early-stopped GD over asymptotic GD such as calibration and a smaller sample complexity.
@@
In the classical finite $d$-dimensional setting, the sample complexity of the empirical risk minimizer (ERM) for logistic regression is well-studied (Ostrovskii & Bach, 2021; Kuchelmeister & van de Geer, 2024; Hsu & Mazumdar, 2024; Chardon et al., 2024), where the minimax rate is known to be $\mathcal{O}(d/n)$. Different from theirs, we focus on an overparameterized regime, where the ERM of logistic regression does not even exist. When specialized to their setting, our Theorem 3.2 recovers the comparable $\tilde{\mathcal{O}}(d/n)$ rate.
@@
The work by Bach (2014) considered one-pass SGD for logistic regression assuming strong convexity around the true model parameter. This strong convexity assumption, however, is prohibitive in our high-dimensional settings.
There is a line of works (Sur & Candès, 2019; Candès & Sur, 2020) focused on the existence of ERM for logistic regression in a propotional limit setting (assuming that $n,d\to\infty$ in a fixed ratio, see also (Chardon et al., 2024) in the finite-dimensional setting). This is quite apart from our focus, where ERM never exists due to overparameterization.
@@
There are logistic risk bounds of early-stopped GD (and one-pass SGD) developed in the noiseless cases, assuming a separable population distribution (Ji & Telgarsky, 2018; Shamir, 2021; Telgarsky, 2022; Schliserman & Koren, 2024). These results do not imply any benefits of early stopping, as their setting is noiseless. In comparison, we consider overparameterized logistic regression with a strictly non-separable population distribution, where the risk of overfitting is prominent. In this case, our results suggest that early stopping plays a significant role in preventing overfitting.
@@
In the boosting literature, an early work by Zhang & Yu (2005) showed that boosting methods (that can be interpreted as coordinate descent) with early stopping are consistent in the classification sense; related refined studies for boosting with the squared loss with early stopping were also provided by Bühlmann & Yu (2003). The paper is also notable for giving the first proof of boosting methods converging to the maximum margin solution (Zhang & Yu, 2005, Appendix D), which was later refined with rates by (Telgarsky, 2013). Their results can be converted to GD. In particular, related concepts were used to prove consistency of early-stopped GD for shallow networks in the lazy regime (Ji et al., 2021). In contrast with the present work that focuses on high-dimensional cases, the preceding works only deal with finite-dimensional settings. Moreover, none of those works provide lower bounds for interpolating estimators and tight links to the regularization path which are provided in the present work.